find the slope or grade of the treadmill when the height is 0.4 ft and lenght is 5 ft
Answer by jim_thompson5910(35256)

Slope = Rise/Run

Slope = 0.4/5

Slope = 0.08

To find the grade, simply multiply by 100 to convert it to a percentage. So this treadmill is at a 100*0.08 = 8% grade